# Mergers all the way down: stellar collisions and kinematics of a dense hierarchically forming massive star cluster in a dwarf starburst

### v1

@@ Abstract
Recent high-resolution observations indicate that the progenitors of globular clusters (GCs) at high redshifts had high average stellar surface densities above $10^{5}$ M pc-2. Studies of the internal structure and kinematics of the clusters, however, remain out of reach. Numerical simulations are necessary to decipher the origin of the zoo of spatio-kinematic features found in present-day GCs. Here we study star cluster formation in a star-by-star hydrodynamical simulation of a low-metallicity starburst occurring during a merger of two gas-rich dwarf galaxies. The simulation accounts for the multiphase interstellar medium, stellar radiation, winds and supernovae, and the accurate small-scale gravitational dynamics near massive stars. We also include prescriptions for stellar collisions and tidal disruption events by black holes. Gravitationally bound star clusters up to $\sim 2\times 10^{5}$ M form dense with initial half-mass radii of 0.1similar-toabsent0.1$\sim$ 0.1∼ 0.11 pc. The most massive cluster approaches the observed high-redshift surface densities throughout its hierarchical and dissipative assembly. The cluster also hosts a collisionally growing very massive star of $\sim 1000$ M that will eventually collapse, forming an intermediate mass black hole. The assembly leaves an imprint in the spatio-kinematic structure of the cluster. The younger half of stars is more centrally concentrated, rotates faster, and its velocity distribution is more radially biased at outer radii. The older population is more round in shape, rotates slowly, its velocity distribution is isotropic and its velocity dispersion is higher. These results provide a possible explanation for a subset of multiple population features observed in GCs such as NGC 104/47 Tuc.
@@ 1 Introduction
Young massive star clusters are among the densest gravitationally bound stellar structures (McLaughlin & van der Marel, 2005; Portegies Zwart et al., 2010; Misgeld & Hilker, 2011). The most extreme examples are super star clusters in local starburst galaxies such as the M82 (McCrady & Graham, 2007; Levy et al., 2024), and young massive star clusters observed through gravitational lensing in the high-redshift Universe, such as in the Sunrise Arc and the Cosmic Gems (Vanzella et al., 2023; Adamo et al., 2024). In both cases, the highest resolved effective surface densities ($\Sigma_{\mathrm{eff}}$) within the effective radius ($R_{\mathrm{eff}}$; projected half-mass or half-light radius) exceed $10^{5}$ $\mathrm{M}_{\odot}$ pc-2. The smallest $R_{\mathrm{eff}}$ of these massive ($M_{\mathrm{cluster}}$$>10^{5}$ $\mathrm{M}_{\odot}$) clusters are less than 1 pc. Runaway stellar collisions and black hole growth (Gold et al., 1965; Spitzer & Saslaw, 1966; Portegies Zwart et al., 2004) as well as chemical self-enrichment through gas retention (Krause et al., 2016; Vink, 2018) have been suggested to take place at such high densities. However, the details of the internal structure of the densest clusters still remain unresolved.
In nearby systems, where the internal structure of clusters can be resolved down to individual stars, clusters that are both young and massive are rare. The R136 cluster in the Large Magellanic Cloud (LMC) is one of the best examples with an age of $\sim 1.5$ Myr (Crowther et al., 2016). The total mass of the cluster has been estimated as $>2\times 10^{4}$ $\mathrm{M}_{\odot}$ within the central few pc or up to $\sim 10^{5}$ $\mathrm{M}_{\odot}$ in total (Andersen et al., 2009). The $\Sigma_{\mathrm{eff}}$ of R136 is, however, only a few hundred $\mathrm{M}_{\odot}$ pc-2. Other notable young massive clusters in the Local Group are NGC 3603 (Harayama et al., 2008), Westerlund 1 (Lim et al., 2013), Westerlund 2 (Ascenso et al., 2007) and the Arches cluster (Espinoza et al., 2009) which have mostly higher metallicities than R136 ($Z\sim$ $0.5$ $\mathrm{Z}_{\odot}$). None of these clusters reach $\Sigma_{\mathrm{eff}}$ as high as the starburst or high-redshift massive clusters.
@@
The spatial and kinematic differences of the stellar populations in present-day GCs provide additional information of the initial stellar distribution. Multiple populations (MPs) are identified through stellar light-element variations that are thought to originate in the early stages of the GC evolution and may thus carry information of the GC formation process (Gratton et al. 2001; Carretta et al. 2009; Piotto et al. 2015; Pancino et al. 2017; Marino et al. 2019; see e.g. Renzini et al. 2015; Charbonnel 2016; Bastian & Lardo 2018; Gratton et al. 2019; Milone & Marino 2022 for reviews). Dalessandro et al. (2019) found that the light-element enhanced second population (P2) in dynamically young GCs tends to be more centrally concentrated than the primordial, field-like population (P1). Leitinger et al. (2023) confirmed the possibility of centrally concentrated P2 but also found a number of clusters for which P1 is instead centrally concentrated. The majority of their clusters, however, showed fully mixed populations. In addition, Dalessandro et al. (2024) found that P2 has predominantly a higher $V/\sigma$ compared to the P1 in the majority of their analysed clusters. Such a scenario is supported by hydrodynamical simulations that follow the formation of the P2 through gas accretion and/or winds of evolved intermediate mass stars in a pre-existing massive P1 component (Bekki, 2010; McKenzie & Bekki, 2021; Lacchin et al., 2022). Martens et al. (2023) and Leitinger et al. (2025), on the other hand, only found one or two such clusters in their samples of 25 and 30 GCs and found no differences in the kinematics of P1 and P2 in the majority of their GCs. Finally, the P1 and P2 sometimes exhibit different velocity anisotropies (Richer et al., 2013; Bellini et al., 2015). N-body simulations indicate that primordial spatial or kinematic differences may be lost during the evolution of the GC in a galactic tidal field except for the most dynamically young clusters (Decressin et al., 2008), and the rate of this relaxation depends on the initial velocity and anisotropy properties (Pavlík et al., 2024).
@@
A self-consistent numerical approach to the initial properties of the populations of stars in young massive clusters forming in a realistic galactic environment is therefore needed. An increasing variety of numerical simulations is now able to probe a wide range of interstellar environments that give rise to the formation of massive star clusters (Renaud et al., 2015; Ma et al., 2020; Lahén et al., 2020a; Hirai et al., 2021; Dobbs et al., 2022; Grudić et al., 2023; Sameie et al., 2023; Mayer et al., 2024; Calura et al., 2024; Andersson et al., 2024; Gutcke, 2024; Reina-Campos et al., 2025; Williams et al., 2025; Matsui et al., 2025). Both observations and simulations agree that galactic systems exhibiting high interstellar gas pressures and high star formation densities harbour the most massive star clusters (Johnson et al., 2017; Adamo et al., 2020; Lahén et al., 2020a). However, because dense star clusters are inherently collisional systems, models of their formation and internal evolution should account for stellar two and few body interactions (e.g. Aarseth 2003; Heggie & Hut 2003). A number of recent hydrodynamical studies have explored star cluster formation and/or evolution at such high fidelity: Fujii et al. (2021), Grudić et al. (2022), Farias et al. (2024), Fujii et al. (2024), Polak et al. (2024) and Cournoyer-Cloutier et al. (2024) modelled star cluster formation in individual molecular clouds at very high spatial resolution and including models for stellar evolution. Fujii et al. (2024), Polak et al. (2024) and Cournoyer-Cloutier et al. (2024) showed that massive star clusters can exhibit very high densities and star formation efficiencies, in agreement or even exceeding the high-$z$ surface densities, when the clusters are still embedded. Cloud-scale simulations are, however, limited to a low number of clusters. They do not consider the role of the galactic environment and thus the impact of the galactic tidal field, gaseous inflows and outflows and the complex time and space varying distribution of feedback from other nearby star forming regions. Toward this end, Jo et al. (2024) introduced a method for modelling the dynamics of stars in a galaxy-scale hydrodynamical simulation but did not include a realistic stellar mass function or star formation.
@@ 2.1 Hydrodynamical simulations
We use the smoothed particle hydrodynamics (SPH) code sphgal (Hu et al., 2014, 2016, 2017) that is an updated version of the gadget-3 code (Springel, 2005) with improvements to the numerical and algorithmic accuracy of the hydrodynamical methods. The code solves the non-equilibrium cooling and heating processes at low temperatures ($<3\times 10^{4}$ K) using a chemical network (H2, H+, H, CO, C+, O and free electrons) following the methods of Nelson & Langer (1997), Glover & Mac Low (2007) and Glover & Clark (2012). At higher temperatures tabulated metallicity dependent cooling rates from Wiersma et al. (2009) are used.
@@ 2.1.1 Star formation
measured when the particle was turned into a star particle. The mass of the star particle (fiducial gas mass resolution $\sim 4$ $\mathrm{M}_{\odot}$) is sampled into several stars and the last sampled star exceeding the particle mass is taken from other such particles waiting to be sampled within $R_{\mathrm{J}}$. This ensures that very massive stars (VMSs; $\gtrsim 100$ $\mathrm{M}_{\odot}$) can only form in relatively dense star-forming regions as demonstrated in Lahén et al. 2023 and Lahén et al. (2024). The most massive star formed this way reaches $\sim 330$ $\mathrm{M}_{\odot}$, and more massive stars therefore form only through stellar collisions (Section 2.1.4). New stars are placed around the parent particle in a Gaussian position and velocity distribution with standard deviation equal to 0.1 pc and 0.1 km s-1 respectively in each cartesian direction. To prevent the creation of tightly bound binaries, we resample positions and velocities of stars that are bound to another new star within a semimajor axis of less than $5\times 10^{-3}$ pc. We also check at the end of the routine that no star is placed closer than $10^{-3}$ pc to another new star, and shift the positions of such stars by an additional Gaussian 0.01 pc. Random initialisation of such binaries is however negligibly rare.
@@ 2.1.2 Stellar feedback
We consider mass and energy release through stellar winds and supernovae. Events that we consider as instantaneous are core-collapse supernovae (zero age main sequence mass 888840404040 $\mathrm{M}_{\odot}$, Chieffi & Limongi 2004), pair-instability supernovae (107.2107.2107.2107.2203.4203.4203.4203.4 $\mathrm{M}_{\odot}$; according to the exploding helium core masses from Heger & Woosley 2002) and asymptotic giant branch winds (0.50.50.50.58888 $\mathrm{M}_{\odot}$, Karakas 2010). Winds of massive stars ($>9$ $\mathrm{M}_{\odot}$) are treated with a continuous momentum-conserving approach. The wind mass loss rates and velocities depend on the initial stellar mass and age, interpolated using the Bonn Optimized Stellar Tracks (BoOST; stellar masses 9999500500500500 $\mathrm{M}_{\odot}$; Szécsi et al. 2022) at metallicity $Z=$$0.01$ $\mathrm{Z}_{\odot}$. We further describe in Section 2.1.4 our treatment of stellar mergers and stellar masses that exceed the upper mass limit of the tables. Chemical enrichment is followed by tracking the mass of 13 chemical elements per particle: H, He, N, C, O, Si, Al, Na, Mg, Fe, S, Ca and Ne.
@@ 2.1.4 Stellar collisions
To our knowledge, at the time of starting the simulation there were no stellar models exceeding 500similar-toabsent500$\sim$ 500∼ 500600600600600 $\mathrm{M}_{\odot}$ that would have also provided us with the chemical properties (e.g. Al, Na abundances) on the surface of the star throughout the stellar evolution. Our choice for the current study was thus to remain conservative and use the 500 $\mathrm{M}_{\odot}$ BoOST model at $Z=0.01$ $\mathrm{M}_{\odot}$ to evolve any star that exceeds 500 $\mathrm{M}_{\odot}$. This results in a lower limit for mass loss and the stellar radius. Recent extensions to the range of tables with initial stellar masses beyond $\sim 500$ $\mathrm{M}_{\odot}$ include the PARSEC tables that now reach 2000 $\mathrm{M}_{\odot}$ at low metallicities (Costa et al., 2025). For future applications we may consider updating our models to such tables. However, given that the BoOST models are among the only ones that provide enough details of the chemical species related to GC MPs, we may alternatively explore the possibility of using tables such as provided by Costa et al. (2025) to build informed extrapolation techniques to extend the upper limit of 500 $\mathrm{M}_{\odot}$ using the BoOST models.
@@ 2.2 Initial conditions and simulation parameters
We continue the starburst simulation for a total duration of $\sim 7.5$ Myr until the largest stellar collision product exceeds $\sim 1000$ $\mathrm{M}_{\odot}$, i.e. when the most massive star considerably exceeds the upper limit of stellar evolution models ($\mathrm{M}_{\odot}$$500$ $\mathrm{M}_{\odot}$). By this time, the peak star formation rate of the starburst has passed, the most massive cluster has formed most of its stars, and the majority of the star clusters formed in the simulation have become gas free.
Figure 1: The mass-size (left) and size-surface density (right) distributions of the simulated and observed young star clusters. The pink circles show the general young star cluster population in the final snapshot. The squares indicate the evolution of the most massive cluster in $\sim 0.1$ Myr steps starting when it had a bound mass equal to $25\%$ of its final mass (see Fig. 2), coloured according to its mean stellar age from light to dark shade. The comparison data are observed young star clusters ($<10$ Myr) in the LEGUS survey (grey dots; Brown & Gnedin 2021) and various surveys of strongly lensed clusters at redshifts $z\gtrsim 2.4$ (red symbols; Vanzella et al. 2022a, b, 2023; Adamo et al. 2024; Mowla et al. 2024; Messa et al. 2025), for which we have used the reported half-light radii to compute the effective surface densities. The solid lines show the best-fit mass-size relation and the corresponding size-surface density relation from (Brown & Gnedin, 2021). The red line shows the compactness parameter $C_{5}=(M_{\mathrm{cluster}}/10^{5}\,\mathrm{M}_{\odot})/(R_{\mathrm{eff}}/\mathrm{pc})=1$ (Krause et al., 2016). The dotted lines show sizes for constant surface density in the range $10^{3}$$10^{6}$ $\mathrm{M}_{\odot}$ pc-2 from top to bottom (left) and surface densities for a constant mass in the range $10^{3}$$10^{6}$ $\mathrm{M}_{\odot}$ from bottom to top (right) in steps of 1 dex.

### v2

@@ Abstract
Recent observations indicate that the progenitors of globular clusters (GCs) at high redshifts had high average stellar surface densities above $10^{5}$ Mpc-2. The internal structure and kinematics of the clusters, however, remain out of reach. Numerical simulations are necessary to decipher the origin of spatio-kinematic features in present-day GCs. Here we study star cluster formation in a star-by-star hydrodynamical simulation of a low-metallicity starburst in a merger of two gas-rich dwarf galaxies. The simulation accounts for the multiphase interstellar medium, stellar radiation, winds and supernovae, and the accurate small-scale gravitational dynamics near massive stars. We also include prescriptions for stellar collisions and tidal disruption events by black holes. Gravitationally bound star clusters up to $\sim 2\times 10^{5}$ M form dense with initial half-mass radii of $\sim 0.1$–1 pc. The most massive cluster approaches the observed high-redshift surface densities throughout its hierarchical and dissipative assembly. The cluster also hosts a collisionally growing very massive star of $\sim 1000$ M that will eventually collapse, forming an intermediate mass black hole. The assembly leaves an imprint in the spatio-kinematic structure of the cluster. The youngest stars are more centrally concentrated, they show significant bulk rotation and have radially biased velocity components at outer radii. The older population is more round in shape, rotates slowly, its velocity distribution is isotropic and exhibits higher dispersion. If chemically enriched star formation proceeds mainly in the later stages of cluster assembly, these results provide a possible explanation for some of the multiple population features observed in dynamically young GCs.
@@ 1 Introduction
Young massive star clusters are among the densest gravitationally bound stellar structures (McLaughlin & van der Marel, 2005; Portegies Zwart et al., 2010; Misgeld & Hilker, 2011). The most extreme examples are super star clusters in local starburst galaxies such as the M82 (McCrady & Graham, 2007; Levy et al., 2024), and young massive star clusters observed through gravitational lensing in the high-redshift Universe, such as in the Sunrise Arc and the Cosmic Gems (Vanzella et al., 2023; Adamo et al., 2024). In both cases, the highest resolved effective surface densities ($\Sigma_{\mathrm{eff}}$) within the effective radius ($R_{\mathrm{eff}}$; projected half-mass or half-light radius) exceed $10^{5}$ $\mathrm{M}_{\odot}$ pc-2. The smallest $R_{\mathrm{eff}}$ of these massive ($M_{\mathrm{cluster}}$$10^{5}$ $\mathrm{M}_{\odot}$) clusters are less than 1 pc. Runaway stellar collisions and black hole growth (Gold et al., 1965; Spitzer & Saslaw, 1966; Portegies Zwart et al., 2004) as well as chemical self-enrichment through gas retention (Krause et al., 2016; Vink, 2018) have been suggested to take place at such high densities. However, the details of the internal structure of the densest clusters still remain unresolved.
In nearby systems, where the internal structure of clusters can be resolved down to individual stars, clusters that are both young and massive are rare. The R136 cluster in the Large Magellanic Cloud (LMC) is one of the best examples with an age of $\sim 1.5$ Myr (Crowther et al., 2016). The total mass of the cluster has been estimated as $2\times 10^{4}$ $\mathrm{M}_{\odot}$ within the central few pc or up to $\sim 10^{5}$ $\mathrm{M}_{\odot}$ in total (Andersen et al., 2009). The $\Sigma_{\mathrm{eff}}$ of R136 is, however, only a few hundred $\mathrm{M}_{\odot}$ pc-2. Other notable young massive clusters in the Local Group are NGC 3603 (Harayama et al., 2008), Westerlund 1 (Lim et al., 2013), Westerlund 2 (Ascenso et al., 2007) and the Arches cluster (Espinoza et al., 2009) which have mostly higher metallicities than R136 ($Z\sim$ $0.5$ $\mathrm{Z}_{\odot}$). None of these clusters reach $\Sigma_{\mathrm{eff}}$ as high as the starburst or high-redshift massive clusters.
@@
The spatial and kinematic differences of the stellar populations in present-day GCs provide additional information of the initial stellar distribution. Multiple populations (MPs) are identified through stellar light-element variations that are thought to originate in the early stages of the GC evolution and may thus carry information of the GC formation process (Gratton et al. 2001; Carretta et al. 2009; Piotto et al. 2015; Pancino et al. 2017; Marino et al. 2019; Cordoni et al. 2025; see e.g. Renzini et al. 2015; Charbonnel 2016; Bastian & Lardo 2018; Gratton et al. 2019; Milone & Marino 2022 for reviews). Dalessandro et al. (2019) found that the light-element enhanced second population (P2) in dynamically young GCs tends to be more centrally concentrated than the primordial, field-like population (P1). Leitinger et al. (2023) confirmed the possibility of centrally concentrated P2 but also found a number of clusters for which P1 is instead centrally concentrated. The majority of their clusters, however, showed fully mixed populations. In addition, Dalessandro et al. (2024) found that P2 has predominantly a higher $V/\sigma$ compared to the P1 in the majority of their analysed clusters. This scenario is supported by hydrodynamical simulations of the formation of the P2 through gas accretion and/or winds of evolved intermediate mass stars in a pre-existing cluster of P1 stars (Bekki, 2010; McKenzie & Bekki, 2021; Lacchin et al., 2022). Martens et al. (2023) and Leitinger et al. (2025), on the other hand, only found one or two such clusters in their samples of 25 and 30 GCs and found no differences in the kinematics of P1 and P2 in the majority of their GCs. Finally, the P1 and P2 sometimes exhibit different velocity anisotropies (Richer et al., 2013; Bellini et al., 2015; Cordoni et al., 2025). N-body simulations indicate that primordial spatial or kinematic differences may be lost during the evolution of the GC in a galactic tidal field except for the most dynamically young clusters (Decressin et al., 2008), and the rate of this relaxation depends on the initial velocity and anisotropy properties (Pavlík et al., 2024).
@@
A self-consistent numerical approach to the initial properties of the populations of stars in young massive clusters forming in a realistic galactic environment is therefore needed. An increasing variety of numerical simulations is now able to probe a wide range of interstellar environments that give rise to the formation of massive star clusters (Renaud et al., 2015; Ma et al., 2020; Lahén et al., 2020a; Hirai et al., 2021; Dobbs et al., 2022; Grudić et al., 2023; Sameie et al., 2023; Mayer et al., 2025; Calura et al., 2025; Andersson et al., 2024; Gutcke, 2024; Reina-Campos et al., 2025; Williams et al., 2025; Matsui et al., 2025). Both observations and simulations agree that galactic systems exhibiting high interstellar gas pressures and high star formation densities harbour the most massive star clusters (Johnson et al., 2017; Adamo et al., 2020; Lahén et al., 2020a). However, because dense star clusters are inherently collisional systems, models of their formation and internal evolution should account for stellar two and few body interactions (e.g. Aarseth 2003; Heggie & Hut 2003). A number of recent hydrodynamical studies have explored star cluster formation and/or evolution at such high fidelity: Fujii et al. (2021), Grudić et al. (2022), Farias et al. (2024), Fujii et al. (2024), Polak et al. (2024) and Cournoyer-Cloutier et al. (2024) modelled star cluster formation in individual molecular clouds at very high spatial resolution and including models for stellar evolution. Fujii et al. (2024), Polak et al. (2024) and Cournoyer-Cloutier et al. (2024) showed that massive star clusters can exhibit very high densities and star formation efficiencies, in agreement or even exceeding the high-$z$ surface densities, when the clusters are still embedded. Cloud-scale simulations are, however, limited to a low number of clusters. They do not consider the role of the galactic environment and thus the impact of the galactic tidal field, gaseous inflows and outflows and the complex time and space varying distribution of feedback from other nearby star forming regions. Toward this end, Jo et al. (2024) introduced a method for modelling the dynamics of stars in a galaxy-scale hydrodynamical simulation but did not include a realistic stellar mass function or star formation.
@@ 2.1 Hydrodynamical simulations
We use the smoothed particle hydrodynamics (SPH) code sphgal (Hu et al., 2014, 2016, 2017) that is an updated version of the gadget-3 code (Springel, 2005) with improvements to the numerical and algorithmic accuracy of the hydrodynamical methods. The code solves the non-equilibrium cooling and heating processes at low temperatures ($3\times 10^{4}$ K) using a chemical network (H2, H+, H, CO, C+, O and free electrons) following the methods of Nelson & Langer (1997), Glover & Mac Low (2007) and Glover & Clark (2012). At higher temperatures tabulated metallicity dependent cooling rates from Wiersma et al. (2009) are used.
@@ 2.1.1 Star formation
measured when the particle was turned into a star particle. The mass of the star particle (fiducial gas mass resolution $\sim 4$ $\mathrm{M}_{\odot}$) is sampled into several stars and the last sampled star exceeding the particle mass is taken from other such particles waiting to be sampled within $R_{\mathrm{J}}$. This ensures that very massive stars (VMSs; $\gtrsim 100$ $\mathrm{M}_{\odot}$) can only form in relatively dense star-forming regions as demonstrated in Lahén et al. 2023 and Lahén et al. (2024). The most massive star formed this way reaches $\sim 330$ $\mathrm{M}_{\odot}$, and more massive stars therefore form only through stellar collisions (Section 2.1.4). New stars are placed around the parent particle in a Gaussian position and velocity distribution with standard deviation equal to 0.1 pc and 0.1 km s-1 respectively in each cartesian direction. Each star is initialised as a single star, and we do not consider primordial binaries in this study. To prevent the creation of tightly bound initial binaries, we resample positions and velocities of stars that are bound to another new star within a semimajor axis of less than $5\times 10^{-3}$ pc. We also check at the end of the routine that no star is placed closer than $10^{-3}$ pc to another new star, and shift the positions of such stars by an additional Gaussian 0.01 pc. Random initialisation of such binaries is however negligibly rare.
@@ 2.1.2 Stellar feedback
We consider mass and energy release through radiation, stellar winds and supernovae. Events that we consider to occur instantaneously at the end of the stellar lifetime are core-collapse supernovae (zero age main sequence mass 884040 $\mathrm{M}_{\odot}$, Chieffi & Limongi 2004), pair-instability supernovae (107.2107.2203.4203.4 $\mathrm{M}_{\odot}$; according to the exploding helium core masses from Heger & Woosley 2002) and asymptotic giant branch winds (0.50.588 $\mathrm{M}_{\odot}$, Karakas 2010). The stellar lifetimes are estimated according to the Geneva stellar models (Georgy et al., 2013) for low-mass stars (<9<9 $\mathrm{M}_{\odot}$) and using the Bonn Optimized Stellar Tracks (BoOST; Szécsi et al. 2022) for massive stars (99500500 $\mathrm{M}_{\odot}$). Winds of massive stars (>9>9 $\mathrm{M}_{\odot}$) are treated with a continuous momentum-conserving approach. The wind mass loss rates and velocities depend on the initial stellar mass and age, interpolated using the BoOSTmodels at metallicity Z=Z=$0.01$ $\mathrm{Z}_{\odot}$. We further describe in Section 2.1.4 our treatment of stellar mergers and stellar masses that exceed the upper mass limit of the tables. Chemical enrichment is followed by tracking the mass of 13 chemical elements per particle: H, He, N, C, O, Si, Al, Na, Mg, Fe, S, Ca and Ne.
@@ 2.1.4 Stellar collisions
To our knowledge, at the time of starting the simulation there were no stellar models exceeding $\sim 500$$600$ $\mathrm{M}_{\odot}$ that would have also provided us with the chemical properties (e.g. Al, Na abundances) on the surface of the star throughout the stellar evolution. Our choice for the current study was thus to remain conservative and use the 500 $\mathrm{M}_{\odot}$ BoOST model at $Z=0.01$ $\mathrm{M}_{\odot}$ to evolve any star that exceeds 500 $\mathrm{M}_{\odot}$. This results in a lower limit for mass loss and the stellar radius. Recent extensions to the range of tables with initial stellar masses beyond $\sim 500$ $\mathrm{M}_{\odot}$ include the PARSEC tables that now reach 2000 $\mathrm{M}_{\odot}$ at low metallicities (Costa et al., 2025). For future applications we may consider updating our models to such tables. However, given that the BoOST models are among the only ones that provide enough details of the chemical species related to GC MPs, we may alternatively explore the possibility of using tables such as provided by Costa et al. (2025) to build informed extrapolation techniques to extend the upper limit of 500 $\mathrm{M}_{\odot}$ using the BoOST models.
@@ 2.2 Initial conditions and simulation parameters
We continue the starburst simulation for a total duration of $\sim 7.5$ Myr until the largest stellar collision product exceeds $\sim 1000$ $\mathrm{M}_{\odot}$, i.e. when the most massive star considerably exceeds the upper limit of stellar evolution models ($500$ $\mathrm{M}_{\odot}$). By this time, the peak star formation rate of the starburst has passed, the most massive cluster has formed most of its stars, and the majority of the star clusters formed in the simulation have become gas free.
Figure 1: The mass-size (left) and size-surface density (right) distributions of the simulated and observed young star clusters. The pink open circles show the general young star cluster population in the final snapshot. The squares indicate the evolution of the most massive cluster in $\sim 0.1$ Myr steps starting when it had a bound mass equal to $25\%$ of its final mass (see Fig. 2), coloured according to its mean stellar age from light to dark shade. The comparison data are observed young star clusters (<10<10 Myr) in the LEGUS survey (grey hexbins with the shade from light to dark indicating the number of clusters ranging from 1 to 38 clusters on the left and 1 to 28 clusters on the right; Brown & Gnedin 2021) and various surveys of strongly lensed clusters at redshifts $z\gtrsim 2.4$ (filled symbols; Vanzella et al. 2022a, b, 2023; Adamo et al. 2024; Mowla et al. 2024; Messa et al. 2025), for which we have used the reported half-light radii to compute the effective surface densities. The solid line shows the compactness parameter $C_{5}=(M_{\mathrm{cluster}}/10^{5}\,\mathrm{M}_{\odot})/(R_{\mathrm{eff}}/\mathrm{pc})=1$ (Krause et al., 2016). The dotted lines on the left show sizes corresponding to a constant surface density for masses in the range $10^{3}$$10^{6}$ $\mathrm{M}_{\odot}$ pc-2 from left to right and the dashed lines on the right show the surface densities for a constant mass in the range $10^{3}$$10^{6}$ $\mathrm{M}_{\odot}$ from left to right, in steps of 1 dex.